OpenStudy (anonymous):

solve 5x-2z=-5 9x+4z=16

OpenStudy (rulnick):

If you double the first equation (mult everything by 2) and add the two equations, you get 19x = 6. Can you get it from there?

OpenStudy (rulnick):

Or did I lose you?

OpenStudy (rulnick):

You would get x = 6/19 after dividing both sides by 19. Can you get z from this?

OpenStudy (anonymous):

..... lost

OpenStudy (rulnick):

One way to solve a system of two eq'ns and two unknowns is to add or subtract one equation to/from another. And it's always ok to multiply an equation by a number other than zero.

OpenStudy (rulnick):

So if you start with your original equations, then multiply the first one by 2, you get: 10x-4z=-10 9x+4z=16. And watch what happens when we add these ...

OpenStudy (rulnick):

... we get 19x + 0z = 6 or just 19x = 6.

OpenStudy (anonymous):

the original equation... Solve using elimination . x-2y-4z=-10 2x+y+2z=5 3x-3y-2z=1 ?

OpenStudy (anonymous):

I need to find x,y,z

OpenStudy (rulnick):

That one is kind of a pain, but here's how to do it: Solve the first equation for x: x = -10 + 2y + 4z. Next, replace the x in the other two equations by "-10 + 2y + 4z" ... you see where this goes?

OpenStudy (rulnick):

Once you do this, you will be down to two equations and two unknowns (y and z). It's a bunch of steps, but each step is fairly straightforward.

OpenStudy (rulnick):

Going back to your first problem, just remember to solve for z, too. The full solution would be (x,z) = (6/19, 125/38).

OpenStudy (anonymous):

HHHHHUHUUUUUUUUUUUHHHHHH????

OpenStudy (rulnick):

Sorry to have confused you. Which one are you confused about, the first problem or the second one?

OpenStudy (anonymous):

use elimination method

OpenStudy (anonymous):

Iknoooowwww I have to use the elimination method but im stuck ! ):

OpenStudy (rulnick):

If you solve the first equation for x, you get x = -10 + 2y + 4z. Now replace x in the other two equations with the value above. You get ...

OpenStudy (rulnick):

... you get 2(-10+2y+4z) + y + 2z = 5 when you do this for the middle equation. Do you know how to simplify this?

OpenStudy (anonymous):

-20 + 4y +8z + y + 2z = 5 -20 + 5y +10z = 5 +20 +20 5y+10z=25

OpenStudy (rulnick):

Perfection!

OpenStudy (rulnick):

Now you have to do the same thing to the third equation. A pain, right?

OpenStudy (anonymous):

but how am i supposed to get them down into xyz form ? like no fraction or equations ? ??? S:

OpenStudy (anonymous):

0,-3,4